I have a triagle that one leg is 91 inches and another leg is 96 inches, What is the length of the third leg? I really am stuck on this one.
Answer by checkley77(12844)

If this is a right triangle and neither of these legs is the hypotenuse then:
91^2+96^2=c^2
8,281+9,216=c^2
17,497=c^2
c=sqrt17,497
c=132.28 in. is the ans. for the hypotenuse (third leg).